Shift report 20030213 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Feb 13 2003 - 07:50:18 EST

  • Next message: Flemming Videbaek: "RHIC running and budget"
    ------------------------------------------------------------
    Date: Thu Feb 13 07:50:17 2003, Shift: 00:00-08:00
    Supervisor: AIJ, with: IA
    Log book#14 pages 13 - 14
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from 4:00 am to 7:30 am
    2. Data runs :
      7876,7877,7878 with following settings :
    MRS 90 deg. 1050A
    FS 12 deg. 1/20B
    
      Need more 150k trigger 3 events on actual MRS settings and <br>60k trigger 6 events on actual FS settings.
    
    3. Problems with TPC's (especialy T3) due to high background <br>in the first part of the store.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu Feb 13 2003 - 07:56:22 EST